# SemEval-2025 Task 1: AdMIRe - Advancing Multimodal Idiomaticity Representation

### v1

@@ Abstract
Thomas Pickard1, Aline Villavicencio1,2 Maggie Mi1, Wei He2, Dylan Phelps1, Carolina Scarton1  and Marco Idiart3 1 University of Sheffield, UK 2 University of Exeter, UK 3 Federal University of Rio Grande do Sul, Brazil {tmrpickard1, zmi1, drsphelps1, c.scarton}@sheffield.ac.uk {a.villavicencio, w.he}@exeter.ac.uk, marco.idiart@gmail.com
@@ 1 Introduction
Advances in language modeling of such phenomena have been made in recent years (Zeng and Bhat, 2022; Zeng et al., 2023; He et al., 2024a), but large language models (LLMs) which perform well on general benchmarks (even for well-resourced languages such as English) fail to consistently exhibit good understanding of figurative language (Mi et al., 2024; Phelps et al., 2024). This has an impact on their application in natural language processing activities such as sentiment analysis, understanding and inference tasks and machine translation (Yazdani et al., 2015). For example, due to poor automatic translation of an idiom, the Israeli PM appeared to call the winner of Eurovision 2018 a ‘real cow’ instead of a ‘real darling’!111metro.co.uk.
@@
The task presented here builds on previous SemEval tasks exploring the evaluation of compositional models Marelli et al. (2014), paraphrases and interpretation of noun compounds Hendrickx et al. (2013); Butnariu et al. (2009) and idiomaticity detection Madabushi et al. (2022). We incorporate visual (§3.1) and visual-temporal (§3.2) modalities across two subtasks in an effort to promote the construction of higher-quality semantic representations of idioms. Our dataset incorporates items in both English (EN) and Brazilian Portuguese (PT-BR), and we focus on nominal compounds which have interpretations in literal and idiomatic senses which are both plausible and imageable.
@@ 2.1 Target compound selection
The potentially idiomatic noun compounds used in this study were sourced from existing datasets including NCTTI Tayyar Madabushi et al. (2021), FLUTE Chakrabarty et al. (2022) and MAGPIE Haagsma et al. (2020). For the purposes of this task, we filtered out compositional expressions (e.g. olive oil), focusing exclusively on those that exhibit duality — i.e., expressions that can be interpreted either literally or figuratively depending on the context (e.g. silver bullet). Additional items known to the researchers were also added as candidates. The candidate lists covered expressions in both English (EN) and Brazilian Portuguese (PT-BR).
@@ 2.2 Static Images (Subtask A)
In total, we obtained 100 English and 30 Portuguese potentially idiomatic expressions, with literal and idiomatic context sentences and visual scene descriptions for each item.
@@ 2.3 Image Sequences (Subtask B)
(d) The image depicts an orange-colored apple that appears to be decomposing or decaying…
(e) The image depicts a rustic, burlap sack filled with several bright orange apples…
Figure 1: Subtask A data example for bad apple. Images were generated using Midjourney. Captions are displayed partially.
@@
2.4 Image Generation
2.5 Caption Generation
3 Task Description
@@ 3.2 Subtask B: Image Sequences (Next Image Prediction)
For each target expression, the first two images in either the literal or idiomatic sequence are provided, along with a set of four candidate images for completion of the sequence. The candidates consist of the intended completion, the two associated alternatives and the completion for the other sense (literal/idiomatic) of the NC.
The task is to correctly identify the intended completion image, while also determining whether the depicted sense of the nominal compound (NC) is idiomatic or literal. Examples are shown in Figure 2.
@@ 3.3.1 Subtask A
For subtask A, we set an expected rank ordering of the 5 images which depends on the sense in which the expression is used in the context sentence. The image strongly associated with the target sense should be ranked first, followed by the mildly associated one. The images for the other sense follow, while the ‘distractor’ image is always expected to be least relevant. For instance, if bad apple is used idiomatically in the context sentence then the expected ranking for the images would be: strongly figurative, miildly figurative, mildly literal, strongly literal, distractor. For the images in Figure 1, this would produce $[a,b,e,d,c]$.
@@
Because our expected order of images is somewhat arbitrary (for a literal instance of a given expression, the idiomatic depictions are essentially no more relevant than the distractor), we adopt a weighting of $[3,1,0,0,0]$ for the five image positions; this allows the metric to capture some of the relevant semantics beyond the top image accuracy without penalising systems which permute the order of the low-relevance images. The maximum DCG score obtainable is therefore 3.631.
@@ 3.3.2 Subtask B
Team Test Set Metrics Extended Evaluation Metrics
Rank Top 1 Acc DCG Score Rank Top 1 Acc DCG Score
PALI-NLP 1 0.93 3.52 1 0.83 3.43
dutir914 2 0.93 3.46 3 0.79 3.28
AlexUNLP 3 0.93 3.45 5 0.72 3.22
AIMA 4 0.87 3.44 10 0.48 2.90
daalft 5 0.87 3.43 2 0.81 3.35
PoliTo 6 0.87 3.381 4 0.475 3.20
UCSC NLP T6 7 0.87 3.36 - - -
Zhoumou 8 0.73 3.20 6 0.69 3.21
HiTZ-Ixa 9 0.73 3.13 7 0.58 3.00
TübingenCL 10 0.67 3.16 - - -
Howard University-AI4PC 11 0.67 3.13 - - -
UoR-NCL 12 0.67 3.10 8 0.57 2.96
FJWU_SemEvalSquad 13 0.60 2.90 11 0.47 2.85
JNLP 14 0.53 3.14 9 0.55 3.13
Modgenix 15 0.53 2.82 - - -
YNU-HPCC 16 0.47 2.85 - - -
UMUTeam 17 0.40 2.68 12 0.24 2.52
Table 2: Leaderboard results - Subtask A, English, Text & Images
Team Test Set Metrics Extended Evaluation Metrics
Rank Top 1 Acc DCG Score Rank Top 1 Acc DCG Score
CTYUN-AI 1 0.87 3.51 1 0.64 3.10
daalft 2 0.67 3.07 4 0.33 2.61
JNLP 3 0.67 3.04 3 0.51 2.86
Transformer25 4 0.47 2.82 2 0.54 3.04
ChuenSumi 5 0.40 2.89 5 0.29 2.68
Table 3: Leaderboard results - Subtask A, English, Text Only
4 Participating Systems and Results
@@ 4.1 Results
HiTZ-Ixa 1 1.00 3.51 7 0.45 2.82
dutir914 2 0.92 3.43 2 0.69 3.06
Zhoumou 3 0.85 3.33 3 0.67 3.10
daalft 4 0.77 3.31 4 0.56 2.95
PALI-NLP 5 0.69 3.21 1 0.76 3.23
AlexUNLP 6 0.62 3.09 6 0.51 2.91
UoR-NCL 7 0.54 3.05 5 0.56 2.90
YNU-HPCC 8 0.38 2.92 - - -
UMUTeam 9 0.38 2.57 8 0.18 2.33
HowardUniversity-AI4PC 10 0.23 2.64 - - -
Modgenix 11 0.23 2.57 - - -
@@
daalft 1 0.60 1.00 2 0.23 0.77
PALI-NLP 2 0.60 0.80 1 0.93 1.00
Modgenix 3 0.60 0.60 - - -
@@
daalft 1 1.00 1.00 2 0.60 0.77
PALI-NLP 2 0.80 0.60 1 0.60 0.90
@@ 4.2 Popular Approaches
Participating teams employed a variety of approaches to the AdMIRe subtasks. All teams opted to use prompting methods with large, generative language and vision-language models and/or to fine-tune smaller pretrained models on the task. Popular model families included GPT-4 OpenAI et al. (2024); QWEN Bai et al. (2023); SBERT Reimers and Gurevych (2019); RoBERTa Liu et al. (2019); CLIP Radford et al. (2021) and its variants such as CLIP-ViLT Wang et al. (2024), AltCLIP Chen et al. (2022) and ALIGN Jia et al. (2021), and BLIP Li et al. (2022). There was also some exploration of the recent DeepSeek model family DeepSeek-AI et al. (2025).
@@
Several teams reported that they took steps to try to ensure that the outputs of generative LLMs were useful, including by prompting to constrain generation and by post-processing to parse the text. Three teams working with LLMs observed a bias in these models towards treating potentially idiomatic expressions as idiomatic regardless of the context in which they occur. Other work has described similar challenges when using LLMs for idiomaticity detection Phelps et al. (2024); Mi et al. (2024); He et al. (2024b).
@@
Seven teams describe some kind of data augmentation step, including paraphrasing and backtranslation to generate variations for finetuning, the addition of information about the target compound (by distillation from LLMs) and image variations. Two teams (FJWU_SemEvalSquad and dutir914) generated additional training instances automatically using generative model pipelines. We hope that these datasets will be made public by the authors in support of future research efforts.
@@ 5 Human Evaluation
In order to provide a comparison point with the model-driven systems, we presented subtask A (using the English extended evaluation dataset) to human annotators. 12 (self-described) fluent speakers of English were recruited from among the staff and postgraduate research students of a UK University, and were compensated with a voucher to the value of GBP15 for their time. The annotations were collected using a survey deployed on the Qualtrics online platform [^7] https://www.qualtrics.com/. For each item, annotators were asked whether they are familiar with its idiomatic meaning. If they answered ‘Yes’, they were then asked to use drag-and-drop to rank the candidate images according to how well they represent the meaning of the expression as it is used in the context sentence.
@@ 7 Conclusion
The datasets for both subtasks are small in the context of contemporary machine learning data, which limited how much they could be used to train or fine-tune models. However, we believe that the manual curation of target items, context sentences, image prompts and generated images by native speakers means that the data quality is high.
@@
The AdMIRe dataset contains items in only two languages, both of which are at least moderately well-resourced. Expanding the dataset to more, and more diverse, languages could enable better evaluation of idiomaticity representation.
@@
Similarly, the datasets are likely to reflect biases and limitations present in the tools we used to construct them, especially the image generation and captioning models. While we made efforts to introduce diversity in the people depicted, we encountered some challenges in this regard. For instance, we were unable to identify prompts which would successfully generate images depicting a literal one-armed bandit [^8] Idiomatically, a slot machine.. Expressions like blue blood also presented difficulty, and we note that the depictions of female characters in the dataset tend to be ‘conventionally attractive’ and have somewhat exaggerated facial features.

### v2

@@ Abstract
Thomas Pickard1, Aline Villavicencio1,2 Maggie Mi1, Wei He2, Dylan Phelps1and Marco Idiart3 1 University of Sheffield, UK 2 University of Exeter, UK 3 Federal University of Rio Grande do Sul, Brazil {tmrpickard1, zmi1, drsphelps1}@sheffield.ac.uk {a.villavicencio, w.he}@exeter.ac.uk, marco.idiart@gmail.com
@@ 1 Introduction
Advances in language modeling of such phenomena have been made in recent years (Zeng and Bhat, 2022; Zeng et al., 2023; He et al., 2024a), but large language models (LLMs) which perform well on general benchmarks (even for well-resourced languages such as English) fail to consistently exhibit good understanding of figurative language (Mi et al., 2024; Phelps et al., 2024). This has an impact on their application in natural language processing activities such as sentiment analysis (Williams et al., 2015; Spasić et al., 2020), understanding and inference tasks and machine translation (Yazdani et al., 2015; Syahrir and Hartina, 2021). For example, due to poor automatic translation of an idiom, the Israeli PM appeared to call the winner of Eurovision 2018 a ‘real cow’ instead of a ‘real darling’!111metro.co.uk.
@@
The task presented here builds on previous SemEval tasks exploring the evaluation of compositional models Marelli et al. (2014), paraphrases and interpretation of noun compounds Hendrickx et al. (2013); Butnariu et al. (2009) and idiomaticity detection Tayyar Madabushi et al. (2022). We incorporate visual (§3.1) and visual-temporal (§3.2) modalities across two subtasks in an effort to promote the construction of higher-quality semantic representations of idioms. Our dataset incorporates items in both English (EN) and Brazilian Portuguese (PT-BR), and we focus on nominal compounds which have interpretations in literal and idiomatic senses which are both plausible and imageable.
@@ 2.1 Target compound selection
The potentially idiomatic noun compounds used in this study were sourced from existing datasets including NCTTI Tayyar Madabushi et al. (2021), FLUTE Chakrabarty et al. (2022) and MAGPIE Haagsma et al. (2020), or identified by the researchers during the project. For the purposes of this task, we filtered out compositional expressions (e.g. olive oil), focusing exclusively on those that exhibit duality — i.e., expressions that can be interpreted either literally or figuratively depending on the context (e.g. silver bullet). The candidate lists covered expressions in both English (EN) and Brazilian Portuguese (PT-BR).
@@ 2.2 Static Images (Subtask A)
In total, we obtained 100 English and 55 Portuguese potentially idiomatic expressions, with literal and idiomatic context sentences and visual scene descriptions for each item.
@@
strongly figurative
mildly figurative
distractor
mildly literal
(d) The image depicts a rustic, burlap sack filled with several bright orange apples…
strongly literal
(e) The image depicts an orange-colored apple that appears to be decomposing or decaying…
Figure 1: Subtask A data example for bad apple. Images generated using Midjourney. Captions are displayed partially.
@@ 2.3 Image Sequences (Subtask B)
2.4 Image Generation
2.5 Caption Generation
3 Task Description
@@ 3.2 Subtask B: Image Sequences (Next Image Prediction)
For each target expression, the first two images in either the literal or idiomatic sequence are provided, along with a set of four candidate images for completion of the sequence. The candidates consist of the intended completion, the two associated alternatives and the completion for the other sense (literal/idiomatic) of the NC. The task is to correctly identify the intended completion image, while also determining whether the depicted sense of the nominal compound (NC) is idiomatic or literal. Examples are shown in Figure 2.
@@ 3.3.1 Subtask A
For subtask A, we set an expected rank ordering of the 5 images which depends on the sense in which the expression is used in the context sentence. The image strongly associated with the target sense should be ranked first, followed by the mildly associated one. The images for the other sense follow, while the ‘distractor’ image is always expected to be least relevant. For instance, if bad apple is used idiomatically in the context sentence then the expected ranking for the images would be: strongly figurative, miildly figurative, mildly literal, strongly literal, distractor. For the images in Figure 1, this would produce $[a,b,d,e,c]$.
@@
Because our expected order of images is somewhat arbitrary (for a literal instance of a given expression, the idiomatic depictions are essentially no more relevant than the distractor), after experimentation we adopt a weighting of $[3,1,0,0,0]$ for the five image positions; this allows the metric to capture some of the relevant semantics beyond the top image accuracy without penalising systems which permute the order of the low-relevance images. The maximum DCG score obtainable is therefore 3.631.
Competition rankings for Subtask A are based on top image accuracy, with DCG breaking ties.
@@ B English 3 1
Team Test Set Metrics Extended Evaluation Metrics
Rank Top 1 Acc DCG Score Rank Top 1 Acc DCG Score
ite]cite.PALI-NLPPALI-NLP 1 0.93 3.52 1 0.83 3.43
ite]cite.dutir914dutir914 2 0.93 3.46 3 0.79 3.28
ite]cite.AlexUNLP-NBAlexUNLP-NB 3 0.93 3.45 5 0.72 3.22
ite]cite.AIMAAIMA 4 0.87 3.44 10 0.48 2.90
ite]cite.daalftdaalft 5 0.87 3.43 2 0.81 3.35
ite]cite.PoliToPoliTo 6 0.87 3.381 4 0.475 3.20
ite]cite.UCSC_NLP_T6UCSC NLP T6 7 0.87 3.36 - - -
ite]cite.ZhoumouZhoumou 8 0.73 3.20 6 0.69 3.21
ite]cite.HiTZ-IxaHiTZ-Ixa 9 0.73 3.13 7 0.58 3.00
ite]cite.TueCLTueCL 10 0.67 3.16 - - -
ite]cite.Howard_University-AI4PCHoward University-AI4PC 11 0.67 3.13 - - -
ite]cite.UoR-NCLUoR-NCL 12 0.67 3.10 8 0.57 2.96
ite]cite.FJWU_SquadFJWU_Squad 13 0.60 2.90 11 0.47 2.85
ite]cite.JNLPJNLP 14 0.53 3.14 9 0.55 3.13
ite]cite.ModgenixModgenix 15 0.53 2.82 - - -
ite]cite.YNU-HPCCYNU-HPCC 16 0.47 2.85 - - -
ite]cite.UMUTeamUMUTeam 17 0.40 2.68 12 0.24 2.52
Table 3: Leaderboard results - Subtask A, English, Text & Images
Team Test Set Metrics Extended Evaluation Metrics
Rank Top 1 Acc DCG Score Rank Top 1 Acc DCG Score
ite]cite.CTYUN-AICTYUN-AI 1 0.87 3.51 1 0.64 3.10
ite]cite.daalftdaalft 2 0.67 3.07 4 0.33 2.61
ite]cite.JNLPJNLP 3 0.67 3.04 3 0.51 2.86
ite]cite.Transformer25Transformer25 4 0.47 2.82 2 0.54 3.04
ite]cite.ChuenSumiChuenSumi 5 0.40 2.89 5 0.29 2.68
Table 4: Leaderboard results - Subtask A, English, Text Only
4 Participating Systems and Results
@@ 4.1 Results
ite]cite.HiTZ-IxaHiTZ-Ixa 1 1.00 3.51 7 0.45 2.82
ite]cite.dutir914dutir914 2 0.92 3.43 2 0.69 3.06
ite]cite.ZhoumouZhoumou 3 0.85 3.33 3 0.67 3.10
ite]cite.daalftdaalft 4 0.77 3.31 4 0.56 2.95
ite]cite.PALI-NLPPALI-NLP 5 0.69 3.21 1 0.76 3.23
ite]cite.AlexUNLP-NBAlexUNLP-NB 6 0.62 3.09 6 0.51 2.91
ite]cite.UoR-NCLUoR-NCL 7 0.54 3.05 5 0.56 2.90
ite]cite.YNU-HPCCYNU-HPCC 8 0.38 2.92 - - -
ite]cite.UMUTeamUMUTeam 9 0.38 2.57 8 0.18 2.33
ite]cite.Howard_University-AI4PCHoward University-AI4PC 10 0.23 2.64 - - -
ite]cite.ModgenixModgenix 11 0.23 2.57 - - -
@@
ite]cite.daalftdaalft 1 0.60 1.00 2 0.23 0.77
ite]cite.PALI-NLPPALI-NLP 2 0.60 0.80 1 0.93 1.00
ite]cite.ModgenixModgenix 3 0.60 0.60 - - -
@@
ite]cite.daalftdaalft 1 1.00 1.00 2 0.60 0.77
ite]cite.PALI-NLPPALI-NLP 2 0.80 0.60 1 0.60 0.90
@@ 4.2 Popular Approaches
Participating teams employed a variety of approaches to the AdMIRe subtasks. All teams opted to use prompting methods with large, generative language and vision-language models and/or to fine-tune smaller pretrained models on the task. Popular model families included GPT-4 OpenAI (2024); QWEN Bai et al. (2023); SBERT Reimers and Gurevych (2019); RoBERTa Liu et al. (2019); CLIP Radford et al. (2021) and its variants such as CLIP-ViLT Wang et al. (2024), AltCLIP Chen et al. (2022) and ALIGN Jia et al. (2021), and BLIP Li et al. (2022). There was also some exploration of the recent DeepSeek model family DeepSeek-AI (2025).
@@
Several teams reported that they took steps to try to ensure that the outputs of generative LLMs were useful, including by prompting to constrain generation and by post-processing to parse the text. Three teams working with LLMs observed a bias in these models towards treating potentially idiomatic expressions as idiomatic regardless of the context in which they occur. Other work has described similar challenges when using LLMs for idiomaticity detection Phelps et al. (2024); Mi et al. (2024); He et al. (2024b). See also Khoshtab et al. (2025) for discussion of prompting methods for LLM figurative language processing.
@@
Seven teams describe some kind of data augmentation step, including paraphrasing and backtranslation to generate variations for finetuning, the addition of information about the target compound (by distillation from LLMs) and image variations. Two teams (FJWU_Squad and dutir914) generated additional training instances automatically using generative model pipelines. We hope that these datasets will be made public by the authors in support of future research efforts.
@@ 5 Human Evaluation
In order to provide a comparison point with the model-driven systems, we presented subtask A (using the English extended evaluation dataset) to human annotators. Twelve (self-described) fluent speakers of English were recruited from among the staff and postgraduate research students of a UK University, and were compensated with a voucher to the value of GBP15 for their time. The annotations were collected using a survey deployed on the Qualtrics online platform [^7] https://www.qualtrics.com/. For each item, annotators were asked whether they are familiar with its idiomatic meaning. If they answered ‘Yes’, they were then asked to use drag-and-drop to rank the candidate images according to how well they represent the meaning of the expression as it is used in the context sentence.
@@ 7 Conclusion
The datasets for both subtasks are small in the context of contemporary machine learning data, which limited how much they could be used to train or fine-tune models. However, we believe that the manual curation of target items, context sentences, image prompts and generated images by native speakers means that the data quality is high. Individual idiomatic expressions are encountered infrequently, even in well-resourced languages such as English, yet are generally well-understood by human readers; we consider the AdMIRe dataset to represent a realistic yet achievable challenge for NLP systems.
@@
The AdMIRe dataset contains items in only two languages, both of which are relatively well-resourced. Expanding the dataset to a greater diversity of languages could enable better evaluation of idiomaticity representation.
@@
Similarly, the datasets are likely to reflect biases and limitations present in the tools we used to construct them, especially the image generation and captioning models. While we made efforts to introduce diversity in the people depicted, we encountered some challenges in this regard. For instance, we were unable to identify prompts which would successfully generate images depicting a literal one-armed bandit [^8] Idiomatically, a slot machine.. Expressions like blue blood also presented difficulty, and we note that the depictions of female characters in particular in the dataset tend to be ‘conventionally attractive’ and have somewhat exaggerated facial features.